With a power supply able to produce 125 A of current, how long must the nails stay in the electrolysis chamber to receive 6.49×106 C of charge? Express your answer to three significant figures and include the appropriate units.
Answer #1

The formula can be used  

I x t = C

I = current im amps = 125 A

t = time in seconds = ?

C =  coulombs of charge =  6.49×106 C

t = 6.49×106/125 = 51920 Seconds or 865.33 Minutes or 14.42 Hours

Hence it will take 14.42 hours to receive 6.49×106 C of charge